# Low-rank matrix and tensor approximations: advancing efficiency of machine-learning interatomic potentials

### v1

@@ Abstract
Machine-learning interatomic potentials (MLIPs) have become a mainstay in computationally-guided materials science, surpassing traditional force fields due to their flexible functional form and superior accuracy in reproducing physical properties of materials. This flexibility is achieved through mathematically-rigorous basis sets that describe interatomic interactions within a local atomic environment. The number of parameters in these basis sets influences both the size of the training dataset required and the computational speed of the MLIP. Consequently, compressing MLIPs by reducing the number of parameters is a promising route to more efficient simulations. In this work, we use low-rank matrix and tensor factorizations under fixed-rank constraints to achieve this compression. In addition, we demonstrate that an algorithm with automatic rank augmentation helps to find a deeper local minimum of the fitted potential. The methodology is verified using the Moment Tensor Potential (MTP) model and benchmarked on multi-component systems: a Mo-Nb-Ta-W medium-entropy alloy, molten LiF-NaF-KF, and a glycine molecular crystal. The proposed approach achieves up to 50% compression without any loss of MTP accuracy and can be applied to compress other MLIPs.
@@ I Introduction
Many MLIPs have been developed since 2007. The first of them are Neural Network Potential (NNP) Behler and Parrinello (2007) based on artificial neural networks, Gaussian Approximation Potential (GAP) Bartók et al. (2010) based on Gaussian processes, and the polynomial-like Spectral Neighbor Analysis Potential (SNAP) Thompson et al. (2015) based on spherical harmonics and Moment Tensor Potential (MTP) Shapeev (2016) based on the tensors of inertia of atomistic environments. Most of recently developed MLIPs are based on neural networks Wang et al. (2018); Pun et al. (2019); Batzner et al. (2022); Batatia et al. (2022) and some of them are polynomial-like Drautz (2019a); Hodapp and Shapeev (2024). In Ref. Deringer, Caro, and Csányi (2019); Zhang et al. (2025), success of applying MLIPs to materials science problems and challenges in their future development are discussed.
@@
In this work, our aim is to verify that compressed MLIPs are as accurate as the base (non-compressed) ones. Hence, after compression, we perform benchmarks on three systems, which were previously studied: the Mo-Nb-Ta-W medium-entropy alloy Hodapp and Shapeev (2024), the molten salt mixture LiF-NaF-KF Rybin et al. (2024), and glycine Rybin, Novikov, and Shapeev (2025a). These systems were chosen not only because the datasets are available, but because MLIPs are, in principle, frequently used to model properties of alloys Rosenbrock et al. (2021); Kostiuchenko et al. (2019), molten salts Porter et al. (2022), and molecular crystals. In the former case, particularly to accelerate the search for the most energetically favorable polymorphs Kapil and Engel (2022); Rybin, Novikov, and Shapeev (2025b); Della Pia et al. (2025); Feng et al. (2025). The success of such calculations is highly dependent on the correct energy ranking of the polymorphs Hunnisett et al. (2024). Molecular crystals typically exhibit complex polymorphic energy landscapes, with numerous structures located within an energy window of a few kilojoules per mole Reilly et al. (2016). Consequently, a model should have DFT accuracy to reliably range polymorphs, which is a good test for compressed MLIP.
We test our methods using Moment Tensor Potential (MTP) as an example model, which is frequently used in computational studies (see, e.g., Refs.Hodapp and Shapeev (2024); Rybin et al. (2024); Rybin, Novikov, and Shapeev (2025a)). This MLIP has two types of parameters to fit: linear and radial. The radial parameters are tensors of the fourth order, and we take advantage of tensor-based decompositions to compress these radial parameters. We demonstrate that the algorithms used enable both reduction of the number of MTP parameters and improvement of accuracy of MTP fitting for the three systems considered. Without loss of generality, the established methodology can be used to compress other MLIPs, for example, ACE Drautz (2019b).
@@ II.3 Matrix and tensor factorization of radial parameters
i.e., if $r_{1}=r_{2}=r_{3}=r$ then we have $(N_{T}+N_{f})r^{2}+(N_{T}+N_{b})r$ radial parameters instead of $N_{T}^{2}N_{f}N_{b}$ in the base MTP. The choice of TT-ranks $\{r_{1},r_{2},r_{3}\}$ governs the trade-off between representational flexibility and compactness.
@@ II.4.2 Optimization with a fixed-rank constraint
The tensor case is similar to MF, except that we minimize the loss function $L(\Xi,G^{(1)},G^{(2)},G^{(3)},G^{(4)})$ with respect to all TT-cores $G^{(k)},k=1,\ldots,4$ and linear parameters $\Xi$ using the BFGS algorithm. Gradients of the loss function with respect to each core are the following:
@@ II.4.3 Optimization with rank augmentation
and start from $r=r_{\min}$. After each of the $s$ iterations of the BFGS algorithm applied to minimize the loss function $L(\Xi,U,V)$, we perform the adaptive update steps if $r<r_{\max}$, where we enrich our approximation with a gradient information. The augmented update steps sequentially increase the rank of the model. A similar approach was used, e.g., in Ref.Dolgov and Savostyanov (2014) for multidimensional linear systems. These steps are presented in detail in Algorithm 1. At each step, the model stores only $r(N_{T}^{2}+N_{f}N_{b})$ radial parameters in factors $U$ and $V$. We refer to the MTP obtained with Algorithm  1 as the MFRA MTP.
@@ III.1 Computational details
We applied developed methods to three systems: the Mo-Nb-Ta-W medium-entropy alloy, the molten salt LiF-NaF-KF (FLiNaK) mixture, and different polymorphs of glycine. The training sets were taken from Refs. Hodapp and Shapeev (2024); Rybin et al. (2024); Rybin, Novikov, and Shapeev (2025a). These data sets were calculated using the VASP package Kresse and Furthmüller (1996) with the projector-augmented wave method Kresse and Joubert (1999). The Perdew-Burke-Ernzerhof generalized gradient approximation (PBE) Perdew, Burke, and Ernzerhof (1996) was used as the exchange–correlation functional, and the DFT-D3 method Grimme et al. (2010) was utilized to account for the dispersion forces in the case of FLiNaK and glycine modeling. A brief description of the calculation parameters is given in Tabel 1 and details might be found in the aforementioned works.
@@
We trained an ensemble of five MTPs of each type: base, MF, R-MF, MFRA, and TF. We fixed the rank $r$ for the calculations with MF, R-MF, and TF. For all calculations with MFRA, we took $s=80$ and $r_{min}=\Delta r$ and increased the rank of the potential until we reached a predetermined $r=r_{max}$. We continued further calculations with this rank. We took $R_{\rm cut}=5$ Å and $N_{b}=8$ for all MTPs. The rest of details on choosing the optimal ranks are given in the next subsection.
We used the weights $w_{e}=1$ eV-2, $w_{f}=0.01$ (eV/Å)-2, and $w_{s}=0$ eV-2 for the fitting on the Mo-Nb-Ta-W and FLiNaK training sets, in the case of glycine we also took into account stresses with $w_{s}=0.001$ eV-2, since in the latter case we had to perform structural relaxation as demonstrated below.
@@ III.2 Determining the optimal ranks on the example of Mo-Nb-Ta-W
Here, we discuss the choice of the optimal rank for the MF, R-MF, and TF MTPs. We refer to optimal rank as the rank that yields a reduction (or compression) of not less than 20% of the radial parameters of the base (non-compressed) MTP and at which the loss function takes its minimum value. We demonstrate the results for the ensembles of MTPs of the 14th level trained on Mo-Nb-Ta-W.
@@
We start with the MF and R-MF MTPs. The dependence of the loss function calculated on the MoNaTaW validation set on the rank of the factorized matrix of MTP radial parameters is shown in Figure  2 (a) (MF) and Figure  2 (b) (R-MF). For MF, the loss function significantly decreases up to the rank of 4 and then continues to decrease more gradually to the rank of 8 (which corresponds to 384 radial parameters) and has the minimum value for this rank. The loss function obtained with R-MF demonstrates similar behavior and reaches a minimum for the rank of 6 corresponding to 252 radial parameters. Thus, the optimal ranks for the MF and R-MF MTPs of the 14th level fitted on the Mo-Nb-Ta-W training set are $r^{\star}_{\text{MF}}=8$ and $r^{\star}_{\text{R-MF}}=6$, respectively.
@@
Figure 2 (c) demonstrates the dependence of the loss function obtained using the TF MTP with the $(r_{1},r_{2},r_{3})$ TT-ranks. The number of radial parameters is given in parentheses. The loss function decreases significantly up to $(r_{1},r_{2},r_{3})=(4,4,4)$ and then flattens. The optimal TT-ranks are $(r_{1}^{*},r_{2}^{*},r_{3}^{*})=(4,8,4)$ which corresponds to 304 radial parameters.
@@
We chose the optimal ranks for the MTPs of other levels fitted on Mo-Nb-Ta-W, as well as for the potentials trained on FLiNaK and glycine in a similar manner. From Figures 2 (a, b, c) it can be seen that the loss functions obtained for optimal ranks are close to those obtained for the ranks related to compression of 50%. Therefore, in this work, we also investigate the compression of approximately 50%. Histograms with ranks and the number of parameters for potentials of the 14th and 18th levels trained on Mo-Nb-Ta-W and FLiNaK are given in Figures  3 and  4, respectively. We provide similar histograms for the MTPs of the 12th and 16th levels in supplementary material. We note that we trained MTPs only of the 20th level for glycine and provide the ranks and the number of radial parameters for each MTP in Table  2 together with the fitting errors.
Figure 4: Histogram with the ranks and the number of parameters (linear and radial) for potentials of the 14th and 18th levels fitted on the F-Li-Na-K training set.

### v2

# Low-rank matrix and tensor approximations for compression of machine-learning interatomic potentials

@@ Abstract
Machine-learning interatomic potentials (MLIPs) have become a mainstay in computationally-guided materials science, surpassing traditional force fields due to their flexible functional form and superior accuracy in reproducing physical properties of materials. This flexibility is achieved through mathematically-rigorous basis sets that describe interatomic interactions within a local atomic environment. The number of parameters in these basis sets influences both the size of the training dataset required and the computational speed of the MLIP. Consequently, compressing MLIPs by reducing the number of parameters is a promising route to more efficient simulations. In this work, we use low-rank matrix and tensor factorizations under fixed-rank constraints to achieve this compression. In addition, we demonstrate that an algorithm with automatic rank augmentation helps to find a deeper local minimum of the fitted potential. The methodology is mainly verified using the Moment Tensor Potential (MTP) model and benchmarked on multi-component systems: a Mo-Nb-Ta-W medium-entropy alloy, molten LiF-NaF-KF, and a glycine molecular crystal. The proposed approach achieves up to 50% compression without any loss of MTP accuracy. We also demonstrate that the developed methodology is universal and can be applied to compress other MLIPs on the example of Atomic Cluster Expansion (ACE).
@@ I Introduction
Many MLIPs have been developed since 2007. The first of them are Neural Network Potential (NNP) Behler and Parrinello (2007) based on artificial neural networks, Gaussian Approximation Potential (GAP) Bartók et al. (2010) based on Gaussian processes, and the polynomial-like Spectral Neighbor Analysis Potential (SNAP) Thompson et al. (2015) based on spherical harmonics and Moment Tensor Potential (MTP) Shapeev (2016) based on the tensors of inertia of atomistic environments. Most of recently developed MLIPs are based on neural networks Wang et al. (2018); Pun et al. (2019); Batzner et al. (2022); Batatia et al. (2022) and some of them are polynomial-like Drautz (2019); Hodapp and Shapeev (2024). In Ref. Deringer, Caro, and Csányi (2019); Zhang et al. (2025), success of applying MLIPs to materials science problems and challenges in their future development are discussed.
@@
In this work, our aim is to verify that compressed MLIPs are as accurate as the base (non-compressed) ones. Hence, after compression, we perform benchmarks on three systems, which were previously studied: the Mo-Nb-Ta-W medium-entropy alloy Hodapp and Shapeev (2024), the molten salt mixture LiF-NaF-KF Rybin et al. (2024), and glycine Rybin, Novikov, and Shapeev (2025). These systems were chosen not only because the datasets are available, but because MLIPs are, in principle, frequently used to model properties of alloys Rosenbrock et al. (2021); Kostiuchenko et al. (2019), molten salts Porter et al. (2022), and molecular crystals. In the former case, particularly to accelerate the search for the most energetically favorable polymorphs Kapil and Engel (2022); Rybin, Novikov, and Shapeev (2025); Della Pia et al. (2025); Feng et al. (2025). The success of such calculations is highly dependent on the correct energy ranking of the polymorphs Hunnisett et al. (2024). Molecular crystals typically exhibit complex polymorphic energy landscapes, with numerous structures located within an energy window of a few kilojoules per mole Reilly et al. (2016). Consequently, a model should have DFT accuracy to reliably range polymorphs, which is a good test for compressed MLIP.
We mainly test our methods using Moment Tensor Potential (MTP) as an example model, which is frequently used in computational studies (see, e.g., Refs.Hodapp and Shapeev (2024); Rybin et al. (2024); Rybin, Novikov, and Shapeev (2025)). This MLIP has two types of parameters to fit: linear and radial. The radial parameters are tensors of the fourth order, and we take advantage of tensor-based decompositions to compress these radial parameters. We demonstrate that the algorithms used enable both reduction of the number of MTP parameters and improvement of accuracy of MTP fitting for the three systems considered. Without loss of generality, the established methodology can be used to compress other MLIPs, for example, ACE Drautz (2019). In addition to MTP, we apply the methodology to ACE and on the example of the FLiNaK system we demonstrate that the compressed ACE retains the accuracy of the base ACE.
@@ II.3 Matrix and tensor factorization of radial parameters
i.e., if $r_{1}=r_{2}=r_{3}=r$ then we have $(N_{T}+N_{f})r^{2}+(N_{T}+N_{b})r$ radial parameters instead of $N_{T}^{2}N_{f}N_{b}$ in the base MTP. The choice of TT‐ranks $\{r_{1},r_{2},r_{3}\}$ governs the trade‐off between representational flexibility and compactness.
@@ II.4.2 Optimization with a fixed-rank constraint
The tensor case is similar to MF, except that we minimize the loss function $L(\Xi,G^{(1)},G^{(2)},G^{(3)},G^{(4)})$ with respect to all TT‐cores $G^{(k)},k=1,\ldots,4$ and linear parameters $\Xi$ using the BFGS algorithm. Gradients of the loss function with respect to each core are given in the supplementary material. We also note that for consistency with matrix factorization (MF) and to improve clarity in the tables and figures, we use the notation TF (tensor factorization) instead of TT.
@@ II.4.3 Optimization with rank augmentation
and start from $r=r_{\min}$. After each of the $s$ iterations of the BFGS algorithm applied to minimize the loss function $L(\Xi,U,V)$, we perform the adaptive update steps if $r<r_{\max}$, where we enrich our approximation with a gradient information. The augmented update steps sequentially increase the rank of the model. A similar approach was used, e.g., in Ref.Dolgov and Savostyanov (2014) for multidimensional linear systems. These steps are presented in detail in Algorithm 1. At each step, the model stores only $r(N_{T}^{2}+N_{f}N_{b})$ radial parameters in factors $U$ and $V$. We refer to the MTP obtained with Algorithm 1 as the MFRA MTP.
@@
II.5 Compressed MTP and Equivariant Tensor Network
In Ref.Hodapp and Shapeev (2024), Equivariant Tensor Network (ETN) potential was proposed. An energy $V^{\rm ETN}_{i}\equiv V^{\rm ETN}({\bf\mathfrak{n}}_{i})$ of atomic neighborhood ${\bf\mathfrak{n}}_{i}$ in ETN potential is the following tensor train (TT):
$V^{\rm ETN}_{i}=\sum_{k_{1}}\sum_{k_{2}}\ldots\sum_{k_{d-2}}\sum_{k_{d-1}}\left(\sum_{k_{1}^{\prime}}T^{1}_{k_{1}^{\prime}k_{1}}v_{k_{1}^{\prime}}\right)\left(\sum_{k_{2}^{\prime}}T^{2}_{k_{1}k_{2}^{\prime}k_{2}}v_{k_{2}^{\prime}}\right)\ldots\\ \displaystyle\left(\sum_{k_{d-1}^{\prime}}T^{d-1}_{k_{d-2}k_{d-1}^{\prime}k_{d-1}}v_{k_{d-1}^{\prime}}\right)\left(\sum_{k_{d}^{\prime}}T^{d}_{k_{d-1}k_{d}^{\prime}}v_{k_{d}^{\prime}}\right),$ (9)
where $T^{1}_{k_{1}^{\prime}k_{1}},T^{2}_{k_{1}k_{2}^{\prime}k_{2}},\ldots,T^{d-1}_{k_{d-2}k_{d-1}^{\prime}k_{d-1}},T^{d}_{k_{d-1}k_{d}^{\prime}}$ are the TT cores, and $v_{k_{j}^{\prime}}\equiv v_{k_{j}^{\prime}}({\bf\mathfrak{n}}_{i})$, $j=1,\ldots,d$ are the following feature vectors:
$v_{k}\equiv v_{lmn}=\sum_{j}f_{n}(|r_{ij}|,z_{i},z_{j})Y_{lm}(r_{ij}/|r_{ij}|),$ (10)
where a multi-index $k=(lmn)$ includes the index $l=0,\ldots,L$ of the subspace of the irreducible representation, the dimension $m=-l,-l+1,\ldots,l$ of the subspace, and the number $n=1,\ldots,N(l)$ of the radial channels corresponding to each $l$ (we denote the maximum number of the radial channels by $N_{\rm rad}^{\rm max}$), $f_{n}(|r_{ij}|,z_{i},z_{j})$ is a radial part describing the pairwise interactions, and $Y_{lm}(r_{ij}/|r_{ij}|)$ are spherical harmonics describing an angular part. The $n$-th radial function in the ETN feature vector has the form:
$f_{n}(r_{ij},z_{i},z_{j})=\sum_{\beta\lambda}B_{n\beta\lambda}\varphi^{(\beta)}(|r_{ij}|)A_{\lambda z_{i}z_{j}}.$ (11)
Here, $\varphi^{(\beta)}(|r_{ij}|)$ are polynomial functions, $B_{n\beta\lambda}$ and $A_{\lambda z_{i}z_{j}}$ are sparse tensors to be fitted (see details in Ref. Hodapp and Shapeev (2024)). Thus, in the ETN potential the coefficients in the radial part are the compressed tensors by its construction whereas in the compressed MTP the original tensor of the radial parameters is compressed using the known low-rank approximations.
A sufficient condition to make $V^{\rm ETN}_{i}$ invariant to rotations is to let the TT cores be equivariant mappings of the feature vectors to rotate them correspondingly with a basis change under SO(3). To that end, the Wigner-Eckhart theorem is employed. This theorem states that every tensor with three multi-indices $\{(l_{i},m_{i},n_{i})\}_{i=1,2,3}$ can be represented in the form of equivariant mapping:
$T_{k_{1}k_{2}k_{3}}=\xi_{(l_{1}n_{1})(l_{2}n_{2})(l_{3}n_{3})}C_{(l_{1}m_{1})(l_{2}m_{2})(l_{3}m_{3})},$ (12)
where $\xi_{(l_{1}n_{1})(l_{2}n_{2})(l_{3}n_{3})}$ is the tensor of coefficients to be fitted, and $C_{(l_{1}m_{1})(l_{2}m_{2})(l_{3}m_{3})}$ is the Clebsch-Gordan coefficient that defines the symmetry group.
To improve accuracy of ETN, we can increase a number of TT cores in (9) and a number of parameters in the tensors of the radial coefficients in (11). To obtain an MTP of a reasonable accuracy, we can increase its level and add more basis functions $B_{\alpha}$ in its functional form. This leads to an uncontrolled growth of the number of linear and radial parameters in the base MTP. However, in the compressed MTPs, it is possible to control a number of radial parameters by applying low-rank approximations to the tensor of the radial parameters. Thus, in this work we developed the methods to control the growth of the number of parameters in MTP. We illustrate this fact in the section with the results below.
II.6 Compressed Atomic Cluster Expansion
To demonstrate the applicability of the proposed methodology for compression of other MLIPs, we use it for the compression of the radial part of Atomic Cluster Expansion (ACE) from Ref. Drautz (2019). To construct ACE, a single bond basis function is introduced:
$\phi_{z_{i}z_{j}nlm}(\bm{r}_{ij})=R_{nl}^{z_{i}z_{j}}(|\bm{r}_{ij}|)Y_{lm}(\bm{r}_{ij}/|\bm{r}_{ij}|),$ (13)
where $R_{nl}^{z_{i}z_{j}}(|\bm{r}_{ij}|)$ is a radial part and $Y_{lm}(\bm{r}_{ij}/|\bm{r}_{ij}|)$ are spherical harmonics describing an angular part. The meaning of the indices $l$, $m$, $n$ here is the same as in ETN described above. The single bond atomic basis function has the following form:
$A_{iznlm}=\sum\limits_{j=1}^{N_{\rm nbh}}\delta_{zz_{j}}\phi_{z_{i}z_{j}nlm}(\bm{r}_{ij}).$ (14)
Permutation-invariant many-body basis functions are generated by taking products as follows:
$\bm{A}_{i\bm{z}\bm{n}\bm{l}\bm{m}}=\prod_{t=1}^{\nu}A_{iz^{t}n^{t}l^{t}m^{t}}.$ (15)
The products follow a body order of $\nu+1$. The vectors $z$, $n$, $l$, $m$ have a length of $\nu$. Finally, the energy $E_{i}$ of atom $i$ is:
$E_{i}=\sum\limits_{\bm{z}\bm{n}\bm{l}\bm{m}}\tilde{\bm{c}}_{z_{i}\bm{z}\bm{n}\bm{l}\bm{m}}\bm{A}_{i\bm{z}\bm{n}\bm{l}\bm{m}},$ (16)
where $\tilde{\bm{c}}_{z_{i}\bm{z}\bm{n}\bm{l}\bm{m}}$ are the linear parameters of ACE to be fitted. In this work, we refer to the above ACE as the base ACE.
As for MTP, the radial part of ACE is also an expansion over the polynomial functions $\phi_{k}$:
$R_{nl}^{z_{i}z_{j}}(|\bm{r}_{ij}|)=\sum_{k=1}^{N_{k}}c_{nlk}^{z_{i}z_{j}}\phi_{k}(|\bm{r}_{ij}|),$ (17)
where $c_{nlk}^{z_{i}z_{j}}$ are the radial parameters to be fitted, $N_{k}$ is the number of polynomials. The total number of these parameters is $(N_{T}(N_{T}-1)/2+N_{T})N_{k}N_{\rm rad}^{\rm max}(L+1)$ and this number consists of $(N_{T}(N_{T}-1)/2+N_{T})$ tensors $C^{\rm ACE}$ of the third order including $N_{k}N_{\rm rad}^{\rm max}(L+1)$ parameters. We compressed these tensors of the third order using the matrix factorization. To that end, we reshaped them into a matrix
$\hat{C}^{\rm ACE}=\mathrm{reshape}\bigl(C^{\rm ACE},\;[N_{k},N_{\rm rad}^{\rm max}\cdot(L+1)]\bigr)\;\in\;\mathbb{R}^{N_{k}\times N_{\rm rad}^{\rm max}(L+1)}.$
Similarly to the MF MTP, we then approximate the matrix of the radial parameters $\widehat{C}^{\rm ACE}\;\approx\;U\,V^{\!\top}$ where $U\in\mathbb{R}^{N_{k}\times r}$, $V\in\mathbb{R}^{N_{\rm rad}^{\rm max}(L+1)\times r}$, and $r$ is the chosen rank. The compressed ACE contains only $r(N_{k}+N_{\rm rad}^{\rm max}(L+1))$ radial parameters. To optimize this potential, we use the same algorithm as for the MF MTP.
@@ III.1 Computational details
We applied developed methods to three systems: the Mo-Nb-Ta-W medium-entropy alloy, the molten salt LiF-NaF-KF (FLiNaK) mixture, and different polymorphs of glycine. The training sets were taken from Refs. Hodapp and Shapeev (2024); Rybin et al. (2024); Rybin, Novikov, and Shapeev (2025). These data sets were calculated using the VASP package Kresse and Furthmüller (1996) with the projector-augmented wave method Kresse and Joubert (1999). The Perdew-Burke-Ernzerhof generalized gradient approximation (PBE) Perdew, Burke, and Ernzerhof (1996) was used as the exchange–correlation functional, and the DFT-D3 method Grimme et al. (2010) was utilized to account for the dispersion forces in the case of FLiNaK and glycine modeling. A brief description of the calculation parameters is given in Tabel 1 and details might be found in the aforementioned works. We also note that for the Mo-Nb-Ta-W system we have a validation set of 546 configurations.
@@
We trained an ensemble of five MTPs of each type, namely, base, MF, R-MF, MFRA, and TF to investigate Mo-Nb-Ta-W and FLiNaK, and we trained ensembles of 10 MTPs for the glycine system. We fixed the rank $r$ for the calculations with MF, R-MF, and TF. For all calculations with MFRA, we took $s=80$ and $r_{min}=\Delta r$ and increased the rank of the potential until we reached a predetermined $r=r_{max}$. We continued further calculations with this rank. Practical recommendations for the choice of the hyperparameters of MFRA are given in the supplementary material. We took $N_{b}=8$ for all MTPs. The rest of details on choosing the optimal ranks are given in the next subsection.
For the Mo-Nb-Ta-W system we also trained ensembles of five ETN potentials and compared the dependence of validation errors in predicting energy and forces on the number of parameters for ETN and compressed MTPs. For the FLiNaK system we fitted the base ACE and its compressed version and demonstrated that the matrix factorization can be applied to compression of the radial parameters of ACE. We took ACE with $N_{k}=9$, $N_{\rm rad}^{\rm max}=3$, and $L=2$. All the potentials mentioned in this study were fitted with $R_{\rm cut}=5$ Å.
We used the weights $w_{e}=1$ eV-2, $w_{f}=0.01$ (eV/Å)-2, and $w_{s}=0$ eV-2 for the fitting on the Mo-Nb-Ta-W and FLiNaK training sets, in the case of glycine we also took into account stresses with $w_{s}=0.001$ eV-2, since in the latter case we had to perform structural relaxation as demonstrated below. We utilized the BFGS algorithm for fitting of all the models, the number of BFGS steps was 2000.
Figure 4: Histogram with the ranks and the number of parameters (linear and radial) for potentials of the 14th and 18th levels fitted on the FLiNaK training set.
@@ III.2 Determining the optimal ranks on the example of Mo-Nb-Ta-W
Here, we discuss the choice of the optimal rank for the MF, R-MF, and TF MTPs. We refer to optimal rank as the rank that yields a reduction (or compression) of not less than 20% of the radial parameters of the base (non-compressed) MTP and at which the loss function takes its minimum value. We note that the reduction of 20% of parameters was chosen as a reasonable threshold under which the compression has almost no sense, as the compression of less than 20% of parameters makes this potential similar to the base potential. We demonstrate the results for the ensembles of MTPs of the 14th level trained on Mo-Nb-Ta-W.
@@
We start with the MF and R-MF MTPs. The dependence of the loss function calculated on the Mo-Nb-Ta-W validation set on the rank of the factorized matrix of MTP radial parameters is shown in Figure 2 (a) (MF) and Figure 2 (b) (R-MF). For MF, the loss function significantly decreases up to the rank of 4 and then continues to decrease more gradually to the rank of 8 (which corresponds to 384 radial parameters) and has the minimum value for this rank. The loss function obtained with R-MF demonstrates similar behavior and reaches a minimum for the rank of 6 corresponding to 252 radial parameters. Thus, the optimal ranks for the MF and R-MF MTPs of the 14th level fitted on the Mo-Nb-Ta-W training set are $r=8$ and $r=6$, respectively.
@@
Figure 2 (c) demonstrates the dependence of the loss function obtained using the TF MTP with the $(r_{1},r_{2},r_{3})$ TT-ranks. The number of radial parameters is given in parentheses. The loss function decreases significantly up to $(r_{1},r_{2},r_{3})=(4,4,4)$ and then flattens. The optimal TT-ranks are $(r_{1},r_{2},r_{3})=(4,8,4)$ which corresponds to 304 radial parameters.
@@
We chose the optimal ranks for the MTPs of other levels fitted on Mo-Nb-Ta-W, as well as for the potentials trained on FLiNaK and glycine in a similar manner. From Figures 2 (a, b, c) it can be seen that the loss functions obtained for optimal ranks are close to those obtained for the ranks related to compression of 50%. Therefore, in this work, we also investigate the compression of approximately 50%. Histograms with ranks and the number of parameters for potentials of the 14th and 18th levels trained on Mo-Nb-Ta-W and FLiNaK are given in Figures 3 and 4, respectively. We provide similar histograms for the MTPs of the 12th and 16th levels in the supplementary material. We note that we trained MTPs only of the 20th level for glycine and provide the ranks and the number of radial parameters for each MTP in Table 3 together with the fitting errors.
III.3 Comparison of the compressed MTPs and ETN for Mo-Nb-Ta-W
In Figure 5 the dependence of the energy and force validation errors on the number of parameters of the compressed MTPs of the 14th level and ETN fitted on the Mo-Nb-Ta-W training set is demonstrated. We found that the compressed MTPs require a similar number of parameters as ETN to reach similar energy validation error within the statistical error, but most of them give smaller force validation error than ETN. As opposed to Figure 7 in Ref. Hodapp and Shapeev (2024) in which the base MTP was compared to ETN, the compressed MTPs in Figure 5 do not require more parameters to reach similar accuracy as ETN. Thus, in this work we resolved the problem of growing the number of MTP radial parameters. Our focus in further discussion of the results of this work is on the compressed MTPs and we do not make further comparisons with ETN.